1. Answer the following questions

                                                              →            →            →  [ 1

a) Consider the linear transformation T(x) = proju(x), where u =  0

                                                                                                      3 ] 

Find the matrix for T.



b) Find the matrix for the linear transformation which reflects every vector in R2 across the x-axis and then rotates every vector through an angle of 𝝅/6. (Show all working)



1
Expert's answer
2022-03-29T01:15:51-0400

"a:\\\\T\\left( x \\right) =proj_u\\left( x \\right) =\\frac{\\left( u,x \\right)}{\\left\\| u \\right\\| ^2}u=\\frac{1x_1+0x_2+3x_3}{1^2+0^2+3^2}\\left[ \\begin{array}{c}\t1\\\\\t0\\\\\t3\\\\\\end{array} \\right] =\\\\=\\left[ \\begin{array}{c}\t0.1x_1+0.3x_3\\\\\t0\\\\\t0.3x_1+0.9x_3\\\\\\end{array} \\right] \\\\Matrix: A=\\left[ \\begin{matrix}\t0.1&\t\t0&\t\t0.3\\\\\t0&\t\t0&\t\t0\\\\\t0.3&\t\t0&\t\t0.9\\\\\\end{matrix} \\right] \\\\b:\\\\The\\,\\,reflection: R\\left( x,y \\right) =\\left( x,-y \\right) \\\\The\\,\\,rotation: T\\left( x,y \\right) =\\left( x\\cos \\frac{\\pi}{6}-y\\sin \\frac{\\pi}{6},x\\sin \\frac{\\pi}{6}+y\\cos \\frac{\\pi}{6} \\right) =\\\\=\\left( \\frac{\\sqrt{3}}{2}x-\\frac{1}{2}y,\\frac{1}{2}x+\\frac{\\sqrt{3}}{2}y \\right) \\\\The\\,\\,composition:\\\\T\\left( R\\left( x,y \\right) \\right) =\\left( \\frac{\\sqrt{3}}{2}x-\\frac{1}{2}\\left( -y \\right) ,\\frac{1}{2}x+\\frac{\\sqrt{3}}{2}\\left( -y \\right) \\right) =\\\\=\\left( \\frac{\\sqrt{3}}{2}x+\\frac{1}{2}y,\\frac{1}{2}x-\\frac{\\sqrt{3}}{2}y \\right) \\\\Matrix: A=\\left[ \\begin{matrix}\t\\frac{\\sqrt{3}}{2}&\t\t\\frac{1}{2}\\\\\t\\frac{1}{2}&\t\t-\\frac{\\sqrt{3}}{2}\\\\\\end{matrix} \\right]"